OpenStudy (anonymous):

PLZ HELP?!? graph y=2x+3 and its inverse

OpenStudy (anonymous):

find the y-intercept find the x-intercept them plug the point on the graph and connect the points

OpenStudy (anonymous):

then find the inverse of the function and do the same by finding the y and x intercept to graph it

OpenStudy (anonymous):

ok...sort of following what your saying..

OpenStudy (anonymous):

okay first find the y intercept and tell me what you get

OpenStudy (anonymous):

3? I'm failing and end of semester is Friday....

OpenStudy (anonymous):

that is right now find the x intercept

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

well, no to find the x intercept set y to zero and solve for x

OpenStudy (anonymous):

-1.5?

OpenStudy (anonymous):

yes 0 = 2x+3 -- subtract 3 to both sides -3 = 2x -3/2 = x -1.5 = x

OpenStudy (anonymous):

now you have two points which are (0,3) and (-1.5,0) plot those points on the graph and connect them

OpenStudy (anonymous):

ok what about inverse?

OpenStudy (anonymous):

id just do the same but subtract instead correct?

OpenStudy (anonymous):

you have y = 2x+3 switch the x and y you get x = 2y+3 now solve for y

OpenStudy (anonymous):

so it be the same points but y is -1.5 and x=3?

OpenStudy (anonymous):

not accurately find the inverse first then you will be accurate

OpenStudy (anonymous):

but y is -1.5 right?

OpenStudy (anonymous):

for the inverse yeah

OpenStudy (anonymous):

ok so x is 3?

OpenStudy (anonymous):

no you have y = 2x+3 switch the x and y you get x = 2y+3 - solve for y for inverse function so x = 2y+3 -- subtract 3 to both sides x-3 = 2y -- divide by 2 (x-3)/2 = y-- the inverse now find the intercepts